# MoST: Motion Style Transformer between Diverse Action Contents

### 10 Mar 2024 v1

@@ 1 Introduction
Figure 1: Frequent failure cases in existing methods: (a) A result of MotionPuzzle [13] replicating style motion. (b) A result of Aberman et al. [1] showing complete failure with twisted motion.
@@
The most significant challenge is transfer failure, where the generated motion loses content of the content motion or does not reflect style of the style motion. Transfer failure often occurs when the contents of the style motion and content motion differ from each other, e.g. kick and punch. Certainly, its complexity arises from the significant variation in style expression between different parts of the body for distinct contents. Despite the practical importance of transferring styles between different content, existing approaches have not primarily addressed these cases, nor have they been comprehensively evaluated, either qualitatively or quantitatively. Fig. 1 illustrates representative failure cases in existing methods. They fail to clearly disentangle style from content and appropriately transfer it to the target content. Fig. 1 (a) illustrates the result of MotionPuzzle [13], where the generated motion, intended to be a jump, instead displays a kick. This method tends to replicate the motion pattern of the style motion rather than effectively transferring disentangled style. In Fig. 1 (b), Aberman et al. [1] utterly fails to generate a plausible motion. Aberman et al. recognized this issue, acknowledging that their method is less effective in motions other than walking and running.
@@
To address the primary concern of transfer failure, we design a new framework called MoST and new loss functions for effectively transferring style between different contents. As style manifests in different patterns across various body parts based on content, MoST handles body part-specific style and content features. MoST comprises transformer-based Siamese motion encoders, a part-attentive style modulator (PSM), and a motion generator. In contrast to the previous methods [1, 19, 13] that build two separate networks for encoding content and style features, we introduce Siamese motion encoders capable of simultaneously extracting both features from an input motion. Subsequently, PSM modulates the raw style feature extracted from the style motion to align with the content of the content motion before being inserted into the generator. This modulation enables effective disentanglement of style from the style motion and its expression in the content motion. Furthermore, we propose new loss functions. The style disentanglement loss aims to distinctly separate style and content. This separation enhances the model’s capability to robustly transfer style, irrespective of the content of the style motion.
Our resulting motions are of high quality, eliminating the need for any heuristic post-processing. MoST achieves this by generating global translational motion along with pose sequences. In the encoders and the generator, the global translational feature and body part features exchange their information via attention, ensuring consistency between the generated poses and global translation. Moreover, to mitigate issues such as jittering and foot-skating, we introduce a physics-based loss. We evaluated the proposed model on two representative motion capture datasets [29, 1], verifying its performance both qualitatively and quantitatively. The main contributions of this study are summarized as follows:
@@
  • We propose novel loss functions to improve the model’s ability to disentangle style from content within the motion and maintain style consistency.
@@ 3.4 Part-Attentive Style Modulator
Part-attentive style modulator (PSM) modulates the style feature $S^{S}$ to be more effectively expressed in $M^{C}$. The detailed operation is illustrated in Fig. 2 (b). Initially, $S^{S}$ is expressed within the context of $C^{S}$, and our objective is to derive $\tilde{S}^{S}$, suitable for integration into $C^{C}$. We initiate the process by computing cross-attention between two content features, $C^{S}$ and $C^{C}$, where $C^{(\cdot)}\in\mathbb{R}^{(P+1)\times d}$ is obtained by temporally pooling $Y^{(\cdot)}$. This cross-attention identifies how style should be transmitted from a specific body part of $C^{S}$ to a corresponding body part of $C^{C}$, e.g. the arm in punch to the leg in kick. Subsequently, $\tilde{S}^{S}$ is generated by linear combinations of part-specific style features of $S^{S}\in\mathbb{R}^{(P+1)\times d}$ according to the cross-attention. Consequently, PSM prevents the transmission of motion from an undesired body part of $M^{S}$, i.e. it only transfers style well-disentangled from content in $M^{S}$. As a result, the generated motion preserves the distinct content of $M^{C}$ instead of confused combinations of motions from $M^{S}$ and $M^{C}$. The cross-attention mechanism is applied as
@@ 3.6 Loss
where $\lambda$s denote weights of each regularization term. $R_{vel}$ and $R_{acc}$ are adopted by the method of [22] and [13]. Assuming that $\mathcal{G}(Y^{G},S^{C})$ should be identical to $M^{C}$, it should have the same foot contact timing as $M^{C}$. Therefore, we introduce foot contact regularization as
@@ 4.2 Self Study
In Table 2, we evaluate the impact of the proposed loss functions and PSM. When PSM is applied, CC significantly decreases, confirming that the style modulation prevents the generation of movement in undesired body parts. $L_{D}$ substantially reduces both CC and SC++ , highlighting the importance of disentangling style and content. After achieving better disentanglement, the additional impact of PSM is relatively smaller than before; however, it consistently decrease both metrics.
@@
The effectiveness of loss functions is further demonstrated in the visualization of the style feature space. Fig. 5 (a) illustrates the space of $\tilde{S}^{S}$ before applying $L_{D}$, where the style features have been mixed across certain categories. Fig. 5 (b) shows that $L_{D}$ enforces complete separation among different styles. Furthermore, the transition from Fig. 5 (c) to (b) shows that the raw style features become distinctive through PSM. Fig. 6 shows how PSM modulates $S^{S}$ according to different $M^{C}$. In Fig. 6 (a), each point represents an overlap of every 56 $M^{C}$ instances, as $S^{S}$ is yielded independently of $M^{C}$. After PSM, the data points are spread according to the paired $M^{C}$ in the $\tilde{S}^{S}$ space.
@@
Figure 7: Cross-attention maps of PSM. The highest column (Refer to caption) indicates the body part in the style motion from which the style originates. The highest element (Refer to caption) pinpoints the body part in the content motion that will receive the style. Traj. refers to global translation, and Refer to caption and Refer to caption are indicators related to traj.

### 20 Mar 2024 v2

@@ 1 Introduction
Figure 1: Frequent failure cases in existing methods: (a) A result of MotionPuzzle [14] replicating style motion. (b) A result of Aberman et al. [1] showing complete failure with twisted motion. The character for the visualization is sourced from Mixamo [13].
@@
The most significant challenge is transfer failure, where the generated motion loses content of the content motion or does not reflect style of the style motion. Transfer failure often occurs when the contents of the style motion and content motion differ from each other, e.g. kick and punch. Certainly, its complexity arises from the significant variation in style expression between different parts of the body for distinct contents. Despite the practical importance of transferring styles between different content, existing approaches have not primarily addressed these cases, nor have they been comprehensively evaluated, either qualitatively or quantitatively. Fig. 1 illustrates representative failure cases in existing methods. They fail to clearly disentangle style from source content and appropriately transfer it to the target content. Fig. 1 (a) illustrates the result of MotionPuzzle [14], where the generated motion, intended to be a jump, instead displays a kick. This method tends to replicate the motion pattern of the style motion rather than effectively transferring disentangled style. In Fig. 1 (b), Aberman et al. [1] utterly fails to generate a plausible motion. Aberman et al. recognized this issue, acknowledging that their method is less effective in motions other than walking and running.
@@
To address the primary concern of transfer failure, we design a new framework called MoST and new loss functions for effectively transferring style between different contents. As style manifests in different patterns across various body parts based on content, MoST handles body part-specific style and content features. MoST comprises transformer-based Siamese motion encoders, a part-attentive style modulator (PSM), and a motion generator. In contrast to the previous methods [1, 20, 14] that build two separate networks for encoding content and style features, we introduce Siamese motion encoders capable of simultaneously extracting both features from an input motion. Subsequently, PSM modulates the raw style feature extracted from the style motion to align with the content of the content motion before being inserted into the generator. This modulation enables effective disentanglement of style from the style motion and its expression in the content motion. Furthermore, we propose new loss functions. The style disentanglemfent loss aims to distinctly separate style and content. This separation enhances the model’s robustness in transferring style, irrespective of the content of the style motion.
Our resulting motions are of high quality, eliminating the need for any heuristic post-processing. MoST achieves this by generating global translational motion along with pose sequences. In the proposed framework, the global translational feature and body part features exchange their information via attention, ensuring consistency between the generated poses and global translation. Moreover, to mitigate issues such as jittering and foot-skating, we introduce a physics-based loss. We evaluated the proposed model on two representative motion capture datasets [30, 1], verifying its performance both qualitatively and quantitatively. The main contributions of this study are summarized as follows:
@@
  • We propose novel loss functions to improve the model’s ability to disentangle style from content within the motion and generate plausible motion.
@@ 3.4 Part-Attentive Style Modulator
Part-attentive style modulator (PSM) modulates the style feature $S^{S}$, which originates from $M^{S}$, to be more effectively expressed in $M^{C}$. The detailed operation is illustrated in Fig. 2 (b). Initially, $S^{S}$ is expressed within the context of $C^{S}$, and our objective is to derive $\tilde{S}^{S}$, suitable for integration into $C^{C}$. We initiate the process by computing cross-attention between two content features, $C^{S}$ and $C^{C}$, where $C^{(\cdot)}\in\mathbb{R}^{(P+1)\times d}$ is obtained by temporally pooling $Y^{(\cdot)}$. This cross-attention identifies how style should be transmitted from a specific body part of $C^{S}$ to a corresponding body part of $C^{C}$, e.g. the legs in jump to the arm in punch. Subsequently, $\tilde{S}^{S}$ is generated by linear combinations of part-specific style features of $S^{S}\in\mathbb{R}^{(P+1)\times d}$ according to the cross-attention. Consequently, PSM prevents the transmission of motion from an undesired body part of $M^{S}$, i.e. it only transfers style well-disentangled from content in $M^{S}$. As a result, the generated motion preserves the distinct content of $M^{C}$ instead of confused combinations of motions from $M^{S}$ and $M^{C}$. The cross-attention mechanism is applied as
@@ 3.6 Loss
where $\lambda$s denote weights of each regularization term. $R_{vel}$ and $R_{acc}$ are adopted by the method of [23] and [14]. Assuming that $\mathtt{MoST}(M^{G},M^{C})$ should be identical to $M^{C}$, it should have the same foot contact timing as $M^{C}$. Therefore, we introduce foot contact regularization as
@@ 4.2 Self Study
In Table 2, we evaluate the impact of the proposed loss function and PSM. When PSM is applied, CC significantly decreases, confirming that the style modulation prevents the generation of movement in undesired body parts. $L_{D}$ substantially reduces both CC and SC++ , highlighting the importance of disentangling style and content. The model achieves the best performance when both $L_{D}$ and PSM are employed.
@@
The effectiveness of $L_{D}$ is further demonstrated in the visualization of the style feature space. Fig. 5 (a) illustrates the space of $\tilde{S}^{S}$ before applying $L_{D}$, where the style features have been mixed across certain categories. Fig. 5 (b) shows that $L_{D}$ enforces complete separation among different styles. Furthermore, the transition from Fig. 5 (c) to (b) shows that the raw style features become distinctive through PSM. Fig. 6 shows how PSM modulates $S^{S}$ according to different $M^{C}$. Fig. 6 (a) illustrates the space of $S^{S}$, where each point represents an overlap of every 56 $M^{C}$ instances, as $S^{S}$ is yielded independently of $M^{C}$. After PSM, the data points are spread according to the paired $M^{C}$ in the $\tilde{S}^{S}$ space.
@@
Figure 7: Cross-attention maps of PSM. The highest column (Refer to caption) indicates the body part in the style motion from which the style originates. The highest element (Refer to caption) pinpoints the body part in the content motion that will receive the style. Traj. refers to global translation. The symbols Refer to caption and Refer to caption are indicators related to traj.